OpenStudy (anonymous):

Solve 2x2 - 8x = -7 I need help, this is too confusing, please help

OpenStudy (anonymous):

Is your problem: \[2 * 2 - 8x = -7\] ?

OpenStudy (anonymous):

No, it's 2x^2 - 8x = -7

OpenStudy (anonymous):

Ok, so what you have is a quadratic. You can solve quadratics by making them equal to zero. To do this you will want to move the -7 to the other side to make it equal to zero. So you should have \[2x^2 - 8x + 7 = 0\]

OpenStudy (anonymous):

Alright, this is where I'm confused, I don't know what to do next

OpenStudy (anonymous):

Use the quadratic formula to solve for x http://www.purplemath.com/modules/quads/qform01.gif

OpenStudy (campbell_st):

use the general quadratic formula...

OpenStudy (anonymous):

Where our coefficients (a,b,c) come from \[ax^2+bx+c=0\]

OpenStudy (campbell_st):

or complete the square... its the same process

OpenStudy (anonymous):

Is it \[2\pm \sqrt{2}/2\]

OpenStudy (jdoe0001):

is it? what did you get anyway?

OpenStudy (anonymous):

What do you mean?

OpenStudy (jdoe0001):

well... did you use the quadratic formula as recommended? or did you complete the square?

OpenStudy (anonymous):

completing the square

OpenStudy (anonymous):

@mahone331 Yes, I got the same answer from using the quadratic formula.

OpenStudy (anonymous):

The answer is -60

OpenStudy (anonymous):

While completing the square and solving by factoring are two possible methods, they don't work every time. The quadratic formula will work for imaginary numbers and much larger integers. You should try using the quadratic a lot because you should have it memorized in college algebra.

OpenStudy (anonymous):

Thank you